Find two numbers such that:
1. Their difference is 13 and their product is -40
2. Their sum is 2 and their difference is 128

Answer by stanbon(75887)

Find two numbers such that:
1. Their difference is 13 and their product is -40
x-y = 13
xy = -40
------------
Substitute for "x" and solve for "y":
(y+13)y = -40
y^2 + 13y + 40 = 0
(y-5)(y-8) = 0
If y = 5, x = -8
If y = 8, x = -5
----------------------------
2. Their sum is 2 and their difference is 128
x + y = 2
x - y = 128
-------
2x = 130
x = 65
y = 63
---------
Cheers,
Stan H.
-----------